# Extraction of classical ergotropy

### v1

@@ Abstract
Discussion.– There are a number of issues that limit the applicability of the QA protocol for ergotropy extraction. Let’s consider, for example, an initial state of the harmonic oscillator featuring a central valley, e.g.:
$\rho_{0}(q,p)$ $=\exp{\left[-\frac{(p^{2}+(q-q_{0})^{2})^{2}}{4}+\frac{p^{2}+(q-q_{0})^{2}}{2}\right]}/N\,$, (34)
where $N$ is the normalisation factor, see the sketch in Fig. 1e). Any choice of decrasing function $f$ would result in a Mexican hat shaped Hamiltonian $H_{1}$ which could be very hard to be implemented. Furthermore since any such $H_{1}$ would feature a central peak, not all regions $H_{1}<E$ would be simply connected, unlike those for which $H_{0}<E$, therefore they cannot be smoothly and adiabatically linked to each other. How to get around such problems? In this case one could still apply the QA protocol, but with the physical $H_{1}$ in Eq. (32), which has the same level hypersurfaces as $\rho_{0}$. Having $H_{1}$ the same topology of phase volumes as $H_{0}$, by adiabatically return to $H_{0}$, the state would be mapped onto
$\rho^{\prime}(q,p)$ $=\exp{\left[-\frac{(p^{2}+q^{2})^{2}}{4}+\frac{p^{2}+q^{2}}{2}\right]}/N=\frac{e^{-H^{2}_{0}+H_{0}}}{N}\,$, (35)
which is less energetic than $\rho_{0}$, but not yet passive although it is stationary (because it is a function of $H_{0}$, but not a decreasing one). In fact $\rho^{\prime}$ has a central valley, see the sketch in Fig. 1f). Using the quantum jargoon, one would say that the state $\rho^{\prime}$ features population inversions. In order to decrease the energy further one should permute rings of higher populations with rings of same volume with lower population and higher (average) energy. The smaller the rings thickness, the closer one can get to the ergotropy extraction, by increasing the number of such permutations. Interestingly, physical drivings that indeed achieve such permutations have been described in Refs. [47]. Notably, such permutation protocols defy the conditions of validity of the adiabatic theorem (because they aim at adiabatically connecting topologically inequivalent regions), but, as numerically demonstrated in Ref. [48], they can nonetheless approach the wanted permutation with very high accuracy. Thus the breakdown of the ergodic hypothesis does not seem to be such a dramatic problem for ergotropy extraction (as much as it is not dramatic for the mechanical foundations of thermodynamics [35, 28]).
In general it is expected that the QA protocol works well whenever $H_{0}$ and $\rho_{0}$ have topologically equivalent phase space structure. In particular no problems occur when the regions $H_{0}\leq E$ and $\rho_{0}>\sigma$ are all simply connected as is the case in the first example above.
Conclusions.– By bridging between quantum thermodynamics and classical plasma physics, the present work offers new viewpoints to both fields. On one hand it provides a guiding principle for energy extraction from classical systems based on the QA protocol (or variants thereof, e.g., as discussed above); on the other it suggests that the quantum theory of ergotropy has no genuinely quantum feature. The two pictures, classical and quantum, are fully one the close analogue of the other. Specifically a decomposition into coherent and incoherent part exists just like in the quantum case, corroborating the idea that energy coherences do not constitute a genuine quantum thermodynamic advantage [46].
@@
Acknowledgments.– The author thanks Frank Ernesto Quintela Rodriguez for pointing out the plasma physics literature and for suggesting the relevance of the present results. Discussions with Vasco Cavina and Marcus Bonança are gratefully acknowledged. The material presented here is based on lectures delivered at the University of Pisa, Italy, in the academic year 2024/25.
@@ I Appendix
I.1 Equivalence of Eq. (6) and (20)
Equation (19) holds regardless of the specific choice of decreasing function $f$. Choosing $f=R$, with $R$ in Eq. (1) we get $P_{1}(\Omega)=R^{-1}(E_{1}(\Omega)$). Furthermore with the specific choice $f=R$, the function $\Omega_{1}$ boils down to the identity. In fact we have
$\Omega_{1}(x)$ $=\int d\mathbf{z}\theta(x-R(\rho_{0}(\mathbf{z})))$
$=\int d\mathbf{z}\theta(\rho_{0}(\mathbf{z})-R^{-1}(x))=R(R^{-1}(x))=x\,.$ (36)
Therefore the inverse of $\Omega_{1}$ is also the identity, i.e., $E_{1}(\Omega)=\Omega$, in this specific case. Summing up
$P_{1}(\Omega)=R^{-1}(\Omega)\,$, (37)
showing that Eq. (20) coincides with Eq. (6).
I.2 Proof of Eqs. (27)
We first note that, for any function $f$, it is
$\int d\mathbf{z}\,\rho(\mathbf{z})f(H_{0}(\mathbf{z}))=\int d\mathbf{z}\mathcal{D}[\rho](\mathbf{z}))f(H_{0}(\mathbf{z}))\,.$ (38)
@@
where in the second line we have used the definition (26), we then did the usual changes of variables $\mathbf{z}^{\prime}\to e\to\Phi$ (see, e.g., Eq. 6) and finally used $\int d\Phi\delta(\Phi-y)=1$.

### v2

@@ Abstract
Decomopsition of ergotropy in coherent and incoherent parts
The above unified theory of ergotropy is an indispensable tool for assessing the genuine quantum signatures of ergotropy. Ref. [33] has highlighted that in the quantum case, the ergotropy splits into a coherent and an incoherent part. The standard expectation is that the coherent part reveals the quantum fingerprints of energy extraction in the quantum regime. Here we show that, contrary to that expectation, coherent contributions survive the quantum to classical transition, as the decomposition into a coherent and an incoherent part remains unaltered in the classical case.
The main tool that is needed for this analysis is the classical “dephasing” operator $\mathcal{D}$ relative to the Hamiltonian $H_{0}$
$\mathcal{D}[\rho](\mathbf{z})=\int d\mathbf{z}^{\prime}\,\rho(\mathbf{z}^{\prime})\delta[\Omega_{0}(H_{0}(\mathbf{z}^{\prime}))-\Omega_{0}(H_{0}(\mathbf{z}))]\,.$ (31)
$\mathcal{D}$ homogenises a classical distribution $\rho$ over the eigensurfaces of $H_{0}$, thus rendering it “diagonal” with respect to $H_{0}$ (i.e., $\{\mathcal{D}[\rho],H_{0}\}=0$, where $\{\cdot,\cdot\}$ denotes Poisson brakets) while preserving its energy  [^3] The classical dephasing operator was first introduced in Ref. [58] where it was expressed in the energy parametrisation. Here we express it in the equivalent enclosed-volume representation, which we find more convenient. . Exactly like in the quantum case [33], we have
$\mathcal{E}_{c}[\rho]=\mathcal{E}_{c}^{c}[\rho]+\mathcal{E}_{c}^{i}[\rho]$ (32)
where
$\mathcal{E}_{c}^{c}[\rho]$ $=\beta^{-1}(C[\rho]+D[\mathcal{P}[\mathcal{D}[\rho]]||\rho_{\beta}]-D[\mathcal{P}[\rho]||\rho_{\beta}])$ (33)
$\mathcal{E}_{c}^{i}[\rho]$ $=\mathcal{E}_{c}[\mathcal{D}[\rho]]$ (34)
denote the coherent and incoherent parts, respectively. Here, in full analogy with the quantum case $\mathcal{P}[\rho]$ denotes the passive companion of $\rho$, $\rho_{\beta}=e^{-\beta H_{0}}/Z$ is a thermal state, $D[\cdot||\cdot]$ denotes the Kullback-Leibler divergence, and
$C[\rho]=D[\rho||\mathcal{D}[\rho]]$ (35)
quantifies the “coherence” of $\rho$, that is how much it is distinguisgable from its dephased companion $\mathcal{D}[\rho]$. Equations (32-35) read exactly like their quantum counterparts where $D$ denotes the quantum Kullback-Leibler divergence, and $\mathcal{D}$ is the quantum dephasing operator relative to $\hat{H}_{0}$ [33]. The proof proceeds in a way that is fully analogous to the quantum case, see appendix for details.
Note that both in quantum and classical cases lack of commutation between the state $\rho_{0}$ and the Hamiltonian $H_{0}$, either in the sense of quantum commutators or of Poisson brackets, signals that the state $\rho_{0}$ is not a function of $H_{0}$ hence it is not passive and its ergotropy is not null. Such lack of commutation, in both cases is associated to coherences, which in the classical case can be visualised as inhomogeneities of the phase distributions over the energy eigensurfaces [58, 59].
Discussion
We have developed a unified theory of classical and quantum ergotropy, a.k.a available energy. The theory bridges results obtained in distant parts of the literature, i.e., plasma and astro-physics and quantum thermodynamics, by establishing a common dictionary and unifying them, while greatly extending their realm of applicability.
The theory applies to any system obeying Liouville dynamics, regardless of the dimension $d$ of its phase space. This includes collisionless plasmas ($d=6$), single particle systems in D dimensions (e.g., the harmonic oscillator of the above exameple $d=2$), but also many-body interacting systems ($d=6N$), like interacting gases and self-gravitating matter. Due to the semi-classical limit, we can in fact safely state that the present theory applies at scales ranging from atomic to galactic. We remark however that our derivation of the classical limit relies on the ergodic hypothesis. Further investigations are needed to study the non-ergodic case.
Thanks to the unified picture we were able to address open questions existing both in the quantum and the classical realms. On the classical side, an open pressing question is how to practically realise the time dependent perturbation that does indeed extract the ergotropy. By porting the simple quantum solution across the quantum-classical boundary, we swiftly formulate the formal solution of this open problem. The solution, the QA protocol, is extremely simple and intuitive in hind-sight, but yet never reported so far. We remark that its practical implementation may be greatly hindered, depending on the shape of $\rho_{0}$; e.g., when the eigensurfaces of $H_{0}$ and $\rho_{0}$ are not topologically equivalent, it is impossible to link them smoothly and adiabatically. We leave a discussion of these challenging problems to future studies while anticipating that they could be solved by combining the present solution with known phase-space methods such as those that dynamically swap regions of phase space nested one into the other [60]. Despite these difficulties, our results form the basis for the systematic study of the problem of energy extraction in the classical regime.
On the quantum side an issue that is in the limelight of timely investigation concerns the quantification of genuinely quantum signatures associated to non-equilibrium energy extraction processes. The present unified theory is indispensable for a systematic study of this problem. Here, for example, we have highlighted that coherences (which in classical system are revealed as inhomogeneities of the phase distribution [58]) are a source of ergotropy both in the classical and quantum cases. Furthermore we showed that the recently reported decomposition of ergotropy into coherent and incoherent components, which was suggested to be revealing of the quantum fingerprints of ergotropy [33], survives in the classical regime. This indicates that, contrary to a widespread expectation, energy coherences do not necessarily constitute a genuine quantum thermodynamic advantage, which corroborates the view expressed in Ref. [58]. Most likely, quantum fingerprints should be looked among the genuinely quantum aspects of coherence, e.g., quantum entanglement, as suggested in Ref. [61]. The issue remains open but now we have powerful new tools for further systematic investigations.
Acknowledgments
@@
The author thanks Frank Ernesto Quintela Rodriguez for pointing out the plasma physics literature and Vittorio Giovannetti for pointing out Ref. [34]. Discussions with Vasco Cavina and Marcus Bonança are gratefully acknowledged. The material presented here is based on lectures delivered at the University of Pisa, Italy, in the academic year 2024/25.
@@ I Appendix
To prove Eq. (32) we first note that, for any function $f$, it is
$\int d\mathbf{z}\,\rho(\mathbf{z})f(H_{0}(\mathbf{z}))=\int d\mathbf{z}\mathcal{D}[\rho](\mathbf{z})f(H_{0}(\mathbf{z}))\,.$ (36)
@@
where in the second line we have used the definition (31), we then did the changes of variables $\mathbf{z}^{\prime}\to e\to\Phi$, as, in Eq. Ergotropy of a classical system, and finally used $\int d\Phi\delta(\Phi-y)=1$.